# app.build: A Production Framework for Scaling Agentic Prompt-to-App Generation with Environment Scaffolding

### v1

@@ Abstract
We present app.build (https://github.com/appdotbuild/agent/), an open-source framework that improves LLM-based application generation through systematic validation and structured environments. Our approach combines multi-layered validation pipelines, stack-specific orchestration, and model-agnostic architecture, implemented across three reference stacks. Through evaluation on 30 generation tasks, we demonstrate that comprehensive validation achieves 73.3% viability rate with 30% reaching perfect quality scores, while open-weights models achieve 80.8% of closed-model performance when provided structured environments. The open-source framework has been adopted by the community, with over 3,000 applications generated to date. This work demonstrates that scaling reliable AI agents requires scaling environments, not just models—providing empirical insights and complete reference implementations for production-oriented agent systems.
@@ 1.1 The Production Reliability Gap
While AI coding agents demonstrate impressive capabilities on standard benchmarks of isolated tasks like HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), relying on them to build production-ready applications without human supervision remains infeasible. Recent repository-level systems such as Devin (Labs, 2024) and SWE-agent (Yang et al., 2024) represent significant advances, yet their performance on real-world software engineering tasks reveals a substantial gap between research benchmarks and production requirements.
This gap manifests across multiple dimensions. Function-level benchmarks like HumanEval evaluate isolated code generation but fail to capture system-level concerns including error handling, integration complexity, and production constraints (Liu et al., 2023). Even state-of-the-art systems like AutoCodeRover, achieving 19% efficacy on SWE-bench at \$0.43 per issue (Zhang et al., 2024), demonstrate that raw model capability alone is insufficient for reliable automated software development.
The core challenge lies in treating LLMs as standalone systems rather than components requiring structured environments. Current approaches predominantly focus on making models “smarter” via either training or prompt engineering, but this paradigm fails to address fundamental reliability issues inherent in probabilistic generation. Recent surveys (Jiang et al., 2024; Paul et al., 2024) note the field requires a shift from model-centric to environment-centric design.
@@ 1.2 Our Approach: Environment Scaffolding
Principles.
  1. 1.
    Structured task decomposition. The agent works through an explicit sequence of well-scoped tasks (e.g., schema $\rightarrow$ API $\rightarrow$ UI), each with clear inputs/outputs and acceptance rules.
  2. 2.
    Multi-layered validation. Deterministic checks (linters, type-checkers, unit/smoke tests, runtime logs) run after every significant generation, catching errors early and feeding them back for automatic repair.
  3. 3.
    Runtime isolation. All code executes in isolated sandboxes (containers) with ephemeral state, enabling safe trial-and-error and reproducible re-runs.
  4. 4.
    Model-agnostic integration. The scaffolding is decoupled from any particular LLM; different backends can be swapped without changing the workflow.
Why ES vs. model-centric approaches?
Traditional (model-centric) systems prompt an LLM to generate the full solution in one or few passes, with checks (if any) at the end. ES, in contrast, enforces a guarded, iterative loop: generate $\rightarrow$ validate $\rightarrow$ repair, per sub-task. Figure 1 and Table 1 summarize the contrast.
@@
Table 1: Environment scaffolding (ES) vs. model-centric generation.
Aspect Model-Centric Environment Scaffolding (Ours)
Task decomposition Single/loosely guided multi-step; no fixed structure Explicit pipeline (FSM): schema $\rightarrow$ API $\rightarrow$ UI
Validation Late or ad-hoc checks Integrated per-step: linters, type checks, unit/smoke tests
Error recovery Manual/ad-hoc retries Automatic repair loop using error feedback
Execution isolation Often none; runs on host Isolated containers; reproducible runs
Model dependence Strong (prompt/model specific) Model-agnostic; environment guides behavior
Observability Limited, coarse logs Per-step metrics, artifacts, and logs
@@ 1.3 Contributions
  • Open-Source Framework (app.build). We release an implementation of ES that targets three stacks (TypeScript/tRPC, PHP/Laravel, Python/NiceGUI) and ships with validators and deployment hooks.
  • Empirical Evaluation. Across end-to-end app-building tasks, we quantify the effect of validation layers and iterative repair, and compare multiple LLM backends under the same environment.
@@
  • Methodological Insight. We find that improving the environment (constraints, tests, repair loops) often matters more than scaling the model for production reliability.
  • Community Adoption. The framework has been used to generate thousands of applications in practice, suggesting ES is useful beyond controlled experiments.
2 Background and Related Work
2.1 Agentic Software Engineering
@@
The evolution of AI coding agents has progressed from simple code completion to autonomous software engineering systems capable of repository-level modifications. SWE-bench (Jimenez et al., 2024) established the gold standard for evaluating repository-level understanding with 2,294 real GitHub issues from 12 Python projects. The accompanying SWE-agent (Yang et al., 2024) demonstrated that custom agent-computer interfaces significantly enhance performance, achieving 12.5% pass@1 through careful interface design rather than model improvements.
Repository-level agents have emerged as a distinct research direction. WebArena (Zhou et al., 2024) revealed that even GPT-4 achieves only 14.41% success versus 78.24% human performance in realistic environments, demonstrating that environment design matters more than model capability. GAIA (Mialon et al., 2023) reinforces this with 92% human versus 15% GPT-4 performance on practical tasks. AutoCodeRover (Zhang et al., 2024) combines LLMs with spectrum-based fault localization, achieving 19% efficacy on SWE-bench at \$0.43 per issue. More recently, Agentless (Xia et al., 2024) challenged complex agent architectures with a simple three-phase process (localization, repair, validation) achieving 32% on SWE-bench Lite at \$0.70 cost, suggesting that sophisticated architectures may not always improve performance.
Multi-agent systems have consistently outperformed single-agent approaches. AgentCoder (Huang et al., 2024) employs a three-agent architecture (Programmer, Test Designer, Test Executor) achieving 96.3% pass@1 on HumanEval with GPT-4, compared to 71.3% for single-agent approaches. MapCoder (Islam et al., 2024) extends this with four specialized agents replicating human programming cycles, achieving 93.9% pass@1 on HumanEval and 22.0% on the challenging APPS benchmark. MetaGPT (Hong et al., 2024) demonstrates role-based agents communicating through structured documents, achieving 85.9% pass@1 on HumanEval with 100% task completion on software development tasks.
2.2 Production Quality in Generated Code
Ensuring production-ready AI-generated code requires validation approaches beyond simple correctness testing. Static analysis integration has shown promise, with intelligent code analysis agents combining GPT-3/4 with traditional static analysis to reduce false-positive rates from 85% to 66%. Testing frameworks have evolved to address AI-specific challenges. Test-driven approaches like TiCoder achieve 45.97% absolute improvement in pass@1 accuracy through interactive generation. Property-based testing frameworks show 23.1–37.3% relative improvements over established TDD methods by generating tests that capture semantic properties rather than specific implementations.
AST-based validation provides structural correctness guarantees. AST-T5 leverages Abstract Syntax Trees for structure-aware analysis, outperforming CodeT5 by 2–3 points on various tasks. Industry deployment reveals gaps between offline performance and practical usage. CodeAssist collected 2M completions from 1,200+ users over one year, revealing significant discrepancies between benchmark performance and real-world usage patterns.
2.3 Tree Search
Tree search enhances LLM-based solutions and serves as a way to increase compute budget beyond internal model reasoning token budget. The closest approach is used by Li et al. in S* Scaling (Li et al., 2025) by combining iterative feedback with parallel branches taking different paths toward solving the problem. Sampling more trajectories increases success rate significantly, which is evident by difference in pass@1 and pass@3 often by 30% or more.
2.4 Runtime Isolation and Scaling
Sandboxing is a cornerstone due to web applications requiring much more elaborate testing than running unit tests. It includes setup and teardown of databases and browser emulation. For parallel scaling, we use Dagger.io for its caching capabilities and Docker compatibility.
3 Problem Setup and Method
@@ 3.1 Problem Formulation
Figure 2: app.build architecture expressed through environment scaffolding. The orchestrator plans stages per stack; each sub-task runs in a sandbox, is validated, and only then merged. CI/CD and DB provisioning are integrated.
@@ 3.2 Architecture
High-level design. The app.build agent implements ES with a central orchestrator that decomposes a user’s specification into stack-specific stages and executes each stage inside an isolated sandbox with validation before acceptance. The same workflow applies across supported stacks (TypeScript/tRPC, PHP/Laravel, Python/NiceGUI). Per-stage validators are stack-aware (e.g., ESLint+TypeScript and Playwright for tRPC; PHPStan and feature tests for Laravel; pytest/ruff/pyright for Python), and the platform provisions managed Postgres databases and CI/CD hooks.
@@ 4.2 Prompt Dataset
The evaluation dataset comprises 30 prompts designed to assess system performance across diverse application development scenarios. Independent human contributors with no prior exposure to the app.build system created evaluation prompts. Contributors developed tasks reflecting authentic development workflows from their professional experience. Prompts were filtered to exclude enterprise integrations, AI/ML compute requirements, or capabilities beyond framework scope. Raw prompts underwent automated post-processing using LLMs to anonymize sensitive information and standardize linguistic structure. The resulting dataset consists of 30 prompts spanning a complexity spectrum (low: static/single-page UI; medium: single-entity CRUD; high: multi-entity/custom logic). See the full list of prompts in Appendix A.
4.3 Metrics
@@
  • Validation pass rates by check (AB-01, AB-02, AB-03, AB-04, AB-06, AB-07)
@@ 4.5 Assessor Protocol and Scoring
To systematically assess generated application quality, we implement a structured evaluation protocol comprising six standardized functional checks executed by human assessors. The evaluation reports two independent outcomes: a binary viability indicator ($V$) and a 0–10 quality score ($Q$).
@@
  • AB-03, AB-04, AB-06 (Clickable Sweep): PASS = 1.0, WARN = 0.5, FAIL = 0.0
  • AB-07 (Performance): continuous metric normalized to $[0,1]$
Table 2: Check weights and definitions used in scoring (see rubric in Section 4.5). All checks share equal weight after NA re-normalization; AB-01 and AB-02 are hard gates for Viability $V$.
Check ID Check Description Weight (share) Notes
AB-01 Boot & Home 1/6 Hard gate for Viability $V$
AB-02 Prompt Correspondence 1/6 Hard gate for Viability $V$
@@
AB-06 Clickable Sweep 1/6
AB-07 Performance Metrics 1/6 Continuous; normalized to $[0,1]$
  • Note. See mapping of PASS/WARN/FAIL to numeric scores and viability definition in Section 4.5.
@@ 5 Results
5.1 Environment Scaffolding Impact (TypeScript/tRPC only)

### v2

@@ Abstract
Engineering teams increasingly experiment with LLM agents to synthesize full-stack web applications, yet production reliability and code generation reproducibility remain the blocking issues. Ongoing improvements of foundational models alone do not reliably translate into deployable software; what matters in practice is the environment that constrains, validates, and repairs model outputs.
[…]
This paper frames the problem as a software engineering challenge (reliability, maintainability, and cost in agentic development), provides a reproducible evaluation protocol validated at production scale, and distills lessons for practitioners deploying LLM agents. We release the open-source framework (650+ stars) and an artifact to reproduce the main tables.
Keywords software engineering $\cdot$ code generation $\cdot$ LLM agents $\cdot$ validation $\cdot$ environment scaffolding
@@ 1.1 The Production Reliability Gap
The promise of Large Language Model (LLM) agents for automated software development has attracted significant industrial interest, yet a critical gap persists between research benchmarks and production requirements. While research systems demonstrate impressive capabilities on isolated benchmarks—HumanEval (Chen et al., 2021) leaders achieve 90%+ pass rates on function-level tasks, and LiveCodeBench (Jain et al., 2024) reaches over 80% on real GitHub issues—these metrics do not translate to deployable software in industrial contexts.
The gap manifests in three critical dimensions that affect practitioner adoption:
Reliability under constraints. Production systems must operate within fixed time and cost budgets while maintaining deterministic quality gates. LLMs generate probabilistically, producing syntactically correct code that fails integration tests, violates security policies, or exhibits subtle runtime defects (Liu et al., 2023).
Reproducibility and debugging. When generation fails, practitioners need actionable diagnostics. Model-centric approaches offer little guidance for iterative refinement. Practitioners need structured validation feedback that pinpoints specific failure modes, so that repairs can be targeted effectively.
Economic viability. At scale, token costs and iteration cycles determine feasibility. Closed frontier models like Claude Sonnet 4 (Anthropic, 2024) achieve high success rates but at significant costs. For teams generating hundreds of applications, these costs compound rapidly. The industry needs cost-performance tradeoffs: where can open-weights models substitute for frontier models? What validation overhead is justified by reliability gains?
We claim that leaning on model-only improvements is insufficient. The prevailing approach treats reliability as a model capability problem—scale parameters, improve training data, refine prompts. However, our production experience generating thousands of applications reveals that environment design matters more than model selection for industrial deployment. A frontier model without validation produces unreliable apps; industry needs explicit tradeoffs between cost, speed, and correctness. Production-ready systems require frameworks that integrate validation, isolation, and repair as first-class concerns—not post-hoc additions to model outputs. Recent surveys (Jiang et al., 2024; Paul et al., 2024) note the field requires a shift from model-centric to environment-centric design.
@@ 1.2 Our Approach: Environment Scaffolding
How environment scaffolding works in practice
Environment scaffolding structures LLM-based code generation around four core practices that address production reliability requirements:
Structured task decomposition. Rather than asking the model to generate an entire application at once, we break work into explicit stages (schema $\rightarrow$ Application Programming Interface (API) $\rightarrow$ User Interface (UI)) with defined inputs, outputs, and success criteria. This matches how developers actually build software and makes failures easier to diagnose.
Multi-layered validation. After every generation step, deterministic checks run automatically: linters catch syntax errors, type-checkers verify contracts, unit tests validate logic, and smoke tests ensure the app boots. Failures trigger immediate repair loops before moving forward, preventing error accumulation.
Runtime isolation. Every generation and test runs in an isolated sandbox with ephemeral state. If the model generates code that crashes or corrupts data, the container resets cleanly. This enables aggressive trial-and-error without risk to production systems.
Model-agnostic design. The scaffolding layer sits between your workflow and the LLM, allowing you to swap models (e.g., from Claude to Qwen) without rewriting validation logic. This protects against vendor lock-in and enables cost-performance optimization.
Why this differs from model-centric approaches
[…]
@@
Table 1: Environment Scaffolding (ES) vs. Model-Centric Generation
Aspect Model-Centric ES (Ours)
Task decomp. Single/loosely guided; no fixed structure Explicit FSM: schema $\rightarrow$ API $\rightarrow$ UI
Validation Late or ad-hoc Per-step: linters, types, tests
Error recovery Manual/ad-hoc Auto repair loop w/ feedback
Execution Often on host Isolated containers
Model dep. Strong (prompt-specific) Model-agnostic
Observability Limited logs Per-step metrics, artifacts
@@ 1.3 Contributions
  • Open-Source Framework (app.build). We release an implementation of ES that targets three stacks (TypeScript/tRPC, PHP/Laravel, Python/NiceGUI) and ships with validators and deployment hooks. The framework has gained 650+ GitHub stars and 89 forks, demonstrating practitioner adoption.
  • Two-Tier Empirical Evaluation. We conduct 300 end-to-end generation experiments with automated metrics (success rate, cost, tokens, duration) plus detailed human evaluation on 30 representative prompts with 6-criteria quality rubric. This methodology combines automated consistency checks at scale with manual viability assessment on representative samples.
  • Production-Scale Validation. The framework has been deployed in production since June 2025, generating over 3000 user applications during the first 4 months with hundreds of applications generated daily at peak usage, providing ecological validity beyond controlled experiments.
  • Cost-Performance Analysis. We quantify validation overhead through token usage and cost-per-viable-app metrics, showing open-weights models (Qwen3) achieve 70% success at 8.2x lower cost (\$0.61 vs \$5.01 per viable app), while validation ablations reveal that comprehensive testing increases costs by \$40 per cohort but catches real defects.
@@
  • Methodological Insight. We find that improving the environment (constraints, tests, repair loops) often matters more than scaling the model for production reliability, with lightweight smoke tests and backend validation providing most gains while End-to-End (E2E) browser tests introduce brittleness.
2 Related Work
@@
Repository-level agentic SE (2024-2025). The evolution of AI coding agents has progressed from code completion to autonomous software engineering systems. SWE-bench (Jimenez et al., 2024) established the evaluation standard with 2,294 real GitHub issues from 12 Python projects. Recent agents demonstrate that environment design rivals model capability: OpenHands (Wang et al., 2024), published at ICLR 2025, achieves 53% on SWE-bench Verified through an open platform for generalist agents with agent-computer interfaces. SWE-agent (Yang et al., 2024) showed 12.5% pass@1 through careful interface design rather than model improvements. Contemporary 2024 agents include AutoCodeRover (Zhang et al., 2024), which combines LLMs with spectrum-based fault localization (19% on SWE-bench, \$0.43 per issue), and Agentless (Xia et al., 2024), challenging architectural complexity with a simple three-phase process (localization, repair, validation) achieving 32% on SWE-bench Lite.
Validation and environment scaffolding. Production-ready code generation requires validation beyond correctness testing. While early explorations in this space focused on code change classification (Kniazev, 2008), modern frameworks now integrate validation at multiple layers. Test-driven approaches (Fakhoury et al., 2024) achieve 45.97% absolute improvement in pass@1 through interactive generation with dynamic test feedback. AST-based validation (Gong et al., 2024) provides structural guarantees, with AST-T5 outperforming CodeT5 by 2–3 points through structure-aware pretraining. Tree search methods (Li et al., 2025) demonstrate that scaling compute through iterative refinement and parallel branches can significantly improve success rates. Multi-agent systems (Hong et al., 2024) show that role-based collaboration with structured validation outperforms single-agent approaches, achieving 85.9% pass@1 on HumanEval with 100% task completion on development tasks. For web application generation, sandboxed execution with database provisioning and browser emulation is essential for isolating and validating complex multi-tier systems.
3 Industrial Context & System
@@ 3.1 Problem Formulation
Figure 2: app.build architecture expressed through environment scaffolding. The orchestrator plans stages per stack; each sub-task runs in a sandbox, is validated, and only then merged. Continuous Integration/Continuous Deployment (CI/CD) and database provisioning are integrated.
@@ 3.2 Architecture
High-level design. The app.build agent implements ES with a central orchestrator that decomposes a user’s specification into stack-specific stages and executes each stage inside an isolated sandbox with validation before acceptance. The same workflow applies across supported stacks (TypeScript/tRPC, PHP/Laravel, Python/NiceGUI), selected for their deterministic scaffolding patterns and comprehensive validator availability (TypeScript/ESLint/Playwright, PHPStan/Laravel feature tests, pytest/ruff/pyright). Per-stage validators are stack-aware, and the platform provisions managed Postgres databases and CI/CD hooks.
@@ 4.1 Evaluation Framework
Our evaluation methodology addresses the core challenge of assessing prompt-to-app generation systems: validating that generated applications not only compile and boot, but actually implement the requested functionality correctly. We combine manual-heavy viability assessment with automated checks across 300 end-to-end experiments, where every generated application undergoes deterministic validation followed by human review on a representative subset.
Two-tier evaluation design. We conduct 300 end-to-end generation runs with automated checks measuring boot success, smoke test passage, cost, token usage, and duration. These automated gates provide fast, deterministic feedback but cannot assess functional correctness or usability. For viability assessment, human evaluators systematically test 30 representative applications using a structured 6-criteria rubric (Section 4.4), verifying that generated apps correctly implement requested functionality, handle edge cases, and meet production quality standards. This manual assessment is essential because automated metrics alone cannot detect integration bugs, incorrect business logic, or poor UX that would block production deployment.
Viability vs. quality distinction. We separate binary viability ($V\in\{0,1\}$) from continuous quality ($Q\in[0,10]$). Viability requires only that an application boots successfully and demonstrates basic prompt correspondence—a minimal threshold for deployment consideration. Quality scoring evaluates correctness, completeness, error handling, and code maintainability through systematic human assessment. This distinction reflects industrial practice: automated gates filter non-viable candidates before human review evaluates deployment readiness.
Production-scale validation. Beyond controlled experiments, we validate the framework through production deployment metrics. Since June 2025, the system has generated over 3000 user applications with peak usage exceeding 220 apps/day (Section 5). Production data provides ecological validity showing the framework operates reliably with diverse real-world requirements beyond our curated test set.
Cost-performance tradeoffs. All experiments track token consumption (input/output) and API costs to quantify validation overhead. We report both total cost per generation cohort and cost-per-viable-app to reveal true economics: validation may increase upfront costs but reduces the effective cost of successful outcomes by filtering failures early.
@@ 4.2 Prompt Dataset
The evaluation dataset comprises 30 prompts designed to assess system performance across diverse application development scenarios. Independent human contributors with no prior exposure to the app.build system created evaluation prompts. Contributors developed tasks reflecting authentic development workflows from their professional experience. Prompts were filtered to exclude enterprise integrations, AI/ML compute requirements, or capabilities beyond framework scope. Raw prompts underwent automated post-processing using LLMs to anonymize sensitive information and standardize linguistic structure. The resulting dataset consists of 30 prompts spanning a complexity spectrum (low: static/single-page UI; medium: single-entity Create, Read, Update, Delete (CRUD); high: multi-entity/custom logic). See the full list of prompts in Appendix A.
@@
  • Validation pass rates by check (AB-01 through AB-06, defined in Section 4.4)
@@ 4.4 Assessor Protocol and Scoring
To systematically assess generated application quality, we implement a structured evaluation protocol comprising six standardized functional checks executed by human assessors. The evaluation reports two independent outcomes: a binary viability indicator ($V$) and a 0–10 quality score ($Q$). The complete assessor handbook with detailed grading criteria and example graded applications is publicly available (Kniazev et al., 2025).
@@
  • AB-03 (Create Functionality), AB-04 (View/Edit Operations), AB-05 (Clickable Sweep): PASS = 1.0, WARN = 0.5, FAIL = 0.0
  • AB-06 (Performance): continuous metric normalized to $[0,1]$
Table 2: Check Weights and Definitions Used in Scoring
Check ID Description Weight Notes
AB-01 Boot 1/6 Hard gate for $V$
AB-02 Prompt Correspondence 1/6 Hard gate for $V$
@@
AB-05 Clickable Sweep 1/6
AB-06 Performance 1/6 Normalized to $[0,1]$
  • See Section 4.4 for rubric details. All weights equal after NA re-normalization. AB-01 and AB-02 are hard gates for viability ($V$).
@@ 5 Results
5.1 Production Deployment and Community Adoption
The app.build framework has been deployed in production since June 2025, demonstrating real-world viability beyond controlled experiments. The open-source repository (https://github.com/neondatabase/appdotbuild-agent) has gained significant community traction with 650 stars and 89 forks as of October 2025, indicating strong practitioner interest in environment-first approaches to agentic code generation.
Figure 3: GitHub star growth trajectory for appdotbuild/agent repository showing 13x growth over 5 months (May-October 2025), with inflection point in June 2025 coinciding with production deployment launch. The sustained upward trajectory through October 2025 indicates genuine practitioner adoption rather than transient interest. Data from star-history.com.
Figure 4: Production usage metrics demonstrating real-world deployment scale. Left: Daily application creation and deployment activity showing peak usage of 220+ apps/day in early August 2025. Right: User growth trajectory over 30 days showing rapid adoption spike coinciding with peak usage period, reaching 160+ active users. Data from production database analytics.
Figure 3 shows the repository’s star growth trajectory, revealing an inflection point in June 2025 when the framework reached production maturity. The repository grew from approximately 50 stars to 650+ stars over five months, representing 13x growth with peak velocity exceeding 100 stars per month during August-September 2025. This organic adoption pattern—characterized by sustained acceleration rather than a single viral spike—suggests the framework addresses genuine practitioner needs.
At peak usage, the platform generated hundreds of applications daily (Figure 4, left panel shows peak of 220+ apps/day in early August 2025), with over 3000 user applications generated during the first 4 months (June-October 2025). The concurrent user growth spike (Figure 4, right panel) demonstrates sustained platform adoption beyond initial experimentation. This production-scale validation complements our controlled experiments: while our systematic evaluation uses 30 prompts with detailed human assessment and 300 experiments with automated metrics, the production deployment provides ecological validity showing the framework operates reliably in uncontrolled real-world conditions with diverse user requirements.
The community adoption metrics (650+ stars, 89 forks) position app.build among actively-used open-source agent frameworks, demonstrating that practitioners value systematic environment scaffolding for production reliability over model-only approaches. The correlation between production deployment launch (June 2025) and rapid community growth validates the industrial relevance of our environment-first approach.
5.2 Two-Tier Evaluation Methodology
Our evaluation combines automated checks with manual viability assessment. We conducted 300 end-to-end generation experiments across baseline and ablation conditions, collecting objective metrics (success rate, healthcheck pass rate, cost, duration, token usage) for each run. The automated “healthcheck” corresponds to AB-01 (Boot), while full viability assessment requires both AB-01 and AB-02 (Prompt Correspondence) as defined in Section 4.4. These automated checks enable consistent measurement and cost-effectiveness analysis across configurations. For viability validation, we performed detailed human evaluation on 30 representative prompts using the AB-check rubric (Section 4.4), providing nuanced assessment of functional correctness and production readiness that automated metrics cannot capture.
This two-tier approach reflects industrial practice: automated metrics enable systematic comparison across model architectures and validation strategies, while human evaluation validates that generated applications actually work for their intended purpose. Automated gates filter obviously broken candidates, but human assessors determine whether apps meet production deployment criteria.
5.3 Automated Validation Results at Scale (n=300)
Table 3 presents aggregated results from 300 automated experiments across all conditions. The baseline configuration (Claude Sonnet 4 with full validation) achieved 86.7% automated success rate at \$110.20 total cost for 30 apps. Open-weights models show cost-performance tradeoffs: Qwen3-Coder-480B achieved 70% success at \$12.68 total cost, delivering an 8.2x cost reduction per viable app (\$0.61 vs \$5.01), while validation ablations reveal systematic patterns discussed in subsequent sections.
Table 3: Large-Scale Automated Results Across 300 Experiments
Configuration n Success HC Pass Cost Dur.(s)
Baseline (Claude) 30 86.7% 96.7% \$110.20 478
No Lint 30 93.3% 96.7% \$70.49 496
No Playwright 30 83.3% 93.3% \$86.17 463
No Tests 30 93.3% 100% \$71.05 373
Qwen3-480B 90 70.0% 86.7% \$12.68 629
GPT-OSS-120B 90 30.0% 43.3% \$4.55 628
  • Success = automated healthcheck (AB-01) + template validation (AB-02) passed; excludes template-only apps with zero functionality. HC Pass = healthcheck (AB-01) only; includes both functional apps and non-functional templates. Cost = total for cohort. Dur. = mean per-app duration. Open model experiments used simplified validation pipeline (AB-01 + AB-02 only).
Key findings from automated metrics: (1) Removing comprehensive validation (no_lint, no_tests) increases automated success by +6.7% but reduces costs by \$40, suggesting validators catch real issues at measurable expense. (2) Playwright removal has minimal impact on automated success (-3.3%) while saving \$24, indicating E2E brittleness. (3) Open models achieve viable cost-performance tradeoffs for less critical applications.
5.4 Cost and Token Usage Analysis
Detailed telemetry from 300 experiments reveals systematic resource consumption patterns. The baseline configuration (Claude Sonnet 4, full validation) consumed 27.7M input tokens and 1.8M output tokens across 30 apps, averaging 923K input and 60K output tokens per app. This translates to \$3.67 per app at standard API rates (\$3/M input, \$15/M output).
Table 4: Resource Consumption Breakdown by Configuration
Config In Tok Out Tok Cost Viable
/App /App /App Cost
Baseline 923K 60K \$3.67 \$5.01
No Lint 531K 50K \$2.35 \$2.52
No Playwright 694K 53K \$2.87 \$3.45
No Tests 531K 52K \$2.37 \$2.54
Qwen3-480B 728K 26K \$0.42 \$0.61
GPT-OSS-120B 732K 26K \$0.15 \$0.51
  • Tok/App = tokens per application (K = thousands). Viable Cost = cost per viable app (total cost / viable count). Open models via OpenRouter at reduced rates.
The cost-per-viable-app metric reveals validation overhead: baseline achieves viability at \$5.01 per app (22/30 viable), while removing unit tests reduces this to \$2.54 (24/30 viable) despite similar per-generation costs. This indicates that comprehensive validation both increases initial costs and filters marginal cases, raising the effective cost per successful outcome.
Open-weights models demonstrate dramatic cost advantages: Qwen3-Coder-480B generates viable apps at \$0.61 each (8.2x cheaper than Claude baseline), though at reduced success rates (70% vs 86.7%). For large-scale deployment or less critical applications, this represents a viable engineering tradeoff.
Token efficiency varies by validation configuration: linting and unit tests consume substantial input tokens through multi-round validation cycles (baseline: 923K vs no_tests: 531K), suggesting that validation rigor directly impacts computational cost. The output token counts remain relatively stable (50K-60K), indicating that validation affects iteration count more than generation verbosity.
5.5 Detailed Quality Assessment (Human Evaluation, n=30)